# Variants of the Damascus inequality

### v2

@@ Introduction
with equality if and only if $x_{1}=\cdots=x_{n}$ [2, 4]. Similarly fundamental results—including the Cauchy-Schwarz inequality [11], Jensen’s inequality for convex functions [5], and Chebyshev’s sum inequality for ordered sequences [9]—share common threads of symmetry, convexity, and majorisation that continue to guide modern developments in inequality theory.
@@
Questions of exponent extension as such appear widely in inequality theory, often revealing deeper structural properties and delimiting the scope of particular techniques. A systematic characterisation of all valid exponents not only completes the theoretical picture but also provides practical guidance for applications in optimisation and analysis. For our setting, according to the anonymous referee, the difficulty is reflected by the fact that the function
@@
The case $m=1$ is straightforward (true). When $m=2$, we can show that (1) holds for all values of $n$ using basic calculus argument. We demonstrate this below.
@@
to find the critical points of $f$. Next, we may build up the Sturm’s sequence of the function $2x(x^{4}+1)^{2}f^{\prime}(x)$, using a computational software [^1] One may access Maple or the website https://planetcalc.com/7719/. For more information regarding the Sturm’s sequence, we refer the readers to [1].. We demonstrate this below.
@@ 0 $\infty$
which tells us that $h$ is increasing on $(0,1]$. So h(x)<h(4/5)=305/881h(x)<h(4/5)=-305/881. Also, we claim that $\max\{h(x)\mathrel{\mathop{\ordinarycolon}}x\in[0,+\infty)\}\in(4203/10000,1051/2500)$ and it is attained at some $x\in(223/125,357/200)$. To roughly show this, we first note that $h(x)\geqslant 0$ when $x\geqslant 1$, and $\lim_{x\to+\infty}h(x)=0$. Since $h$ is continuous, we can argue using the extreme value theorem that $h$ has a global maximum. Next, we find critical points of $h$ in $0,+\infty)$. One can verify that $h^{\prime}$ has two nonnegative real roots (e.g. by observing its monotonicity), one at $x=0$, and another at some $x\in(223/125,357/200)$ by the intermediate value theorem. By comparing the images of these points, we prove the claim. Now, let us turn to the case analysis.
@@
We argue that $\Delta_{h}(y)<0$ for $y\in(9/10,14/5)$ by constructing the Sturm’s sequence of $(y^{4}+1)^{3}\Delta_{h}(y)$ and seeing that it has no real roots in this interval. Combining this fact, the continuity of $h$, and $h(1)<0$, we deduce that $h$ is GA-concave on $(9/10,14/5)$ according to Lemma 3.6. This implies
$h(y)+h(z)<2h(\sqrt{yz})=2h\left(\frac{1}{\sqrt{x}}\right)=\frac{2\sqrt{x}(1-x\sqrt{x})}{x^{2}+1}.$
@@
where $\ell(x)\mathrel{\mathop{\ordinarycolon}}=2x^{4}+2x^{3}\sqrt{x}+x^{3}+x^{2}\sqrt{x}+x^{2}-x-\sqrt{x}+1$. One can verify that $\ell$ is positive on $(2/5,4/5)$ (e.g. by making a substitution $t=\sqrt{x}$, then constructing the Sturm’s sequence). Thus $h(x)+h(y)+h(z)<0$ and the proof is complete. $\square$
Now, although $(1,1,1)$ is not a global maximum of $S_{n}^{3}$ under $\mathcal{H}_{3}$ for $n\geqslant 4$, it is the (strict) local maximum. This can be shown by applying the second derivative test to the function $S_{n}^{3}(x,y,1/xy)$. A direct computation shows
@@
to find the critical points of $f$. One can construct the Sturm’s sequence of $2x(x^{2}+1)^{2}f^{\prime}(x)$ to show that it has two positive real roots, one at $x=\sqrt{2}-1$, another at $x=1$. We note that $f(117/500)\leqslant 0$. Combining these facts, we deduce that $f(x)\leqslant 0$ on $117/500,+\infty)$ with equality at $x=1$.
@@
One can check that $\Delta_{h}(w)<0$ for $w\in(127/100,20000000/6145149)$ by using the Sturm’s sequence. Therefore $h$ is GA-concave in this range, and so we deduce
@@
Finally, one can argue using calculus that the above expression is negative for $s$ in the given domain. Thus the proof is complete. $\square$
@@
Having characterised all pairs $(n,m)$ for which (1) holds, we now turn our attention to the non-solutions to this inequality. For convenience, we define the set
@@
For instance, we know that $\mathcal{I}^{2}_{n}=\emptyset$ for all $n$, and $(1/2,3/2,4/3)\in\mathcal{I}^{3}_{n}$ for $n\geqslant 5$. A trivial observation is if $\mathcal{I}_{n}^{m}\neq\emptyset$, then $\mathcal{I}_{n}^{m}$ is a nonempty open subset of $\mathcal{H}_{m}$ Consequently, $\mathcal{I}_{n}^{m}$ has positive (m1)(m-1)-dimensional measure with respect to the induced Lebesgue measure on this surface.
@@
$\lim_{x\to 0^{+}}S_{n}^{3}(x,1/x,1)=\lim_{x\to+\infty}S_{n}^{3}(x,1/x,1)=-1.$
@@
$x\in(0,\rho)\cup(\rho^{\prime},+\infty)\quad\Rightarrow\quad S_{n}^{3}(x,1/x,1)<-\frac{7}{8}.$
@@
$S_{n}^{3}(x,y,z)<-\frac{3}{4},\quad\forall(x,y,z)\in B_{\infty}((\rho,1/\rho,1),\varepsilon_{f}).$ (6)
@@
$\Gamma\mathrel{\mathop{\ordinarycolon}}=\{(x,1/x,1)\mathrel{\mathop{\ordinarycolon}}x\in[\rho,\rho^{\prime}]\}.$
@@
$\{B_{\infty}((x,1/x,1),\varepsilon_{x})\mathrel{\mathop{\ordinarycolon}}(x,1/x,1)\in\Gamma\setminus\{(1,1,1)\}\}\cup\{B_{\infty}((1,1,1)),\varepsilon^{\prime})\}$
@@
$\{B_{\infty}((x_{k},1/x_{k},1),\varepsilon_{k})\}_{k=1}^{N}\subseteq\{B_{\infty}((x,1/x,1),\varepsilon_{x})\}\cup\{B_{\infty}((1,1,1)),\varepsilon^{\prime})\}$
@@
$|x-\tilde{\rho}|\leqslant\varepsilon_{f},\quad|y-1/\tilde{\rho}|\leqslant\varepsilon_{f},\quad|z-1|\leqslant\varepsilon_{f}.$
@@
$\|(x^{\prime},y^{\prime},z^{\prime})-(\rho,1/\rho,1)\|_{\infty}=\|(x,y,z)-(\tilde{\rho},1/\tilde{\rho},1)\|_{\infty}\leqslant\varepsilon_{f}.$
@@
$S_{n}^{3}(x,y,z)=f(x)+f(y)+f(z)<f(x^{\prime})+f(y^{\prime})+f(z^{\prime})=S(x^{\prime},y^{\prime},z^{\prime})<0$
@@
for $x>0$. Assume on the contrary that $\mathcal{I}_{n}^{3}$ is unbounded. Then for every $c>0$, there exists $(x_{c},y_{c},z_{c})\in\mathcal{I}_{n}^{3}$ with $x_{c}<c$ where we assume without loss of generality that $x_{c}\leqslant y_{c}\leqslant z_{c}$. We record two elementary bounds.
@@
In addition to the above, it is natural to ask whether analogous structural properties hold for $\mathcal{I}_{n}^{m}$ when $m$ is arbitrary. Can one prove or disprove that $\mathcal{I}_{n}^{m}$ (i) stays uniformly away from $1$, (ii) is bounded, or (iii) forms an increasing family in $n$? Indeed, the answer to question (i) is negative according to Remark after the proof of Theorem 4.1. We additionally note that the proof of Theorem 5.1 relies crucially on the fact that (1) holds for every $n\in\mathbf{N}$ when $m=2$, a feature that fails for other values of $m$, as shown in the previous sections. Because of this, we may likely think that the answer to question (ii) is negative as well. However, we do not know whether the converse “Theorem 5.2 $\Rightarrow$ Theorem 5.1” is true for any $m$. Moreover, we have not been able to construct an explicit counterexample yet. If it surprisingly turns out to be true, then it would require a new technique to prove. Thus this question, together with question (iii), remains open. Clarifying these properties of $\mathcal{I}_{n}^{m}$ for $m\geqslant 3$ would therefore be an interesting direction for further study, which we leave to the interested readers.

### v3

@@ Introduction
with equality if and only if $x_{1}=\cdots=x_{n}$ [2, 4]. Similarly fundamental results—such as the Cauchy-Schwarz inequality [11], Jensen’s inequality for convex functions [5], and Chebyshev’s sum inequality for ordered sequences [9]—share common threads of symmetry, convexity, and majorisation that continue to guide modern developments in inequality theory.
@@
Questions of exponent extension as such appear widely in inequality theory, often revealing deeper structural properties and delimiting the scope of particular techniques. A systematic characterisation of all valid exponents not only completes the theoretical picture but also provides practical guidance for applications in optimisation and analysis. For our setting, the difficulty is reflected by the fact that the function
@@
The case $m=1$ is straightforward (true). When $m=2$, we can show that (1) holds for all values of $n$ using a basic calculus argument. We demonstrate this below.
@@
to find the critical points of $f$. Next, we may build up Sturm’s sequence of the function $2x(x^{4}+1)^{2}f^{\prime}(x)$, using a computational software [^1] One may access Maple or the website https://planetcalc.com/7719/. For more information regarding Sturm’s sequence, we refer the readers to [1].. We demonstrate this below.
@@ 0 $\infty$
which tells us that $h$ is increasing on $(0,1]$. So h(x)<h(4/5)=305/881h(x)<h(4/5)=-305/881 for $0<x<4/5$. Also, we claim that $\max\{h(x)\mathrel{\mathop{\ordinarycolon}}x\in[0,+\infty)\}\in(4203/10000,1051/2500)$ and it is attained at some $x\in(223/125,357/200)$. To roughly show this, we first note that $h(x)\geqslant 0$ when $x\geqslant 1$, and $\lim_{x\to+\infty}h(x)=0$. Since $h$ is continuous, we can argue using the extreme value theorem that $h$ has a global maximum. Next, we find critical points of $h$ in $0,+\infty)$. One can verify that $h^{\prime}$ has two nonnegative real roots (e.g. by observing its monotonicity), one at $x=0$, and another at some $x\in(223/125,357/200)$ by the intermediate value theorem. By comparing the images of these points, we prove the claim. Now, let us turn to the case analysis.
@@
We may argue that $\Delta_{h}(y)<0$ for $y\in(9/10,14/5)$ by constructing Sturm’s sequence of $(y^{4}+1)^{3}\Delta_{h}(y)$ and seeing that it has no real roots in this interval. Combining this fact, the continuity of $\Delta_{h}(y)$, and $\Delta_{h}(1)<0$, we deduce that $h$ is GA-concave on $(9/10,14/5)$ according to Lemma 3.6. This implies
$h(y)+h(z)\leqslant 2h(\sqrt{yz})=2h\left(\frac{1}{\sqrt{x}}\right)=\frac{2\sqrt{x}(1-x\sqrt{x})}{x^{2}+1}.$
@@
where $\ell(x)\mathrel{\mathop{\ordinarycolon}}=2x^{4}+2x^{3}\sqrt{x}+x^{3}+x^{2}\sqrt{x}+x^{2}-x-\sqrt{x}+1$. One can verify that $\ell$ is positive on $(2/5,4/5)$ (e.g. by making a substitution $t=\sqrt{x}$, then constructing Sturm’s sequence). Thus $h(x)+h(y)+h(z)<0$ and the proof is complete. $\square$
Now, although $(1,1,1)$ is not a global maximum of $S_{n}^{3}$ under $\mathcal{H}_{3}$ for $n\geqslant 4$, it is a (strict) local maximum. This can be shown by applying the second derivative test to the function $S_{n}^{3}(x,y,1/xy)$. A direct computation shows
@@
to find the critical points of $f$. One can construct Sturm’s sequence of $2x(x^{2}+1)^{2}f^{\prime}(x)$ to show that it has two positive real roots, one at $x=\sqrt{2}-1$, another at $x=1$. We note that $f(117/500)\leqslant 0$. Combining these facts, we deduce that $f(x)\leqslant 0$ on $117/500,+\infty)$ with equality at $x=1$.
@@
One can check that $\Delta_{h}(w)<0$ for $w\in(127/100,20000000/6145149)$ by using Sturm’s sequence. Therefore $h$ is GA-concave in this range, and so we deduce that
@@
Finally, one can argue using calculus or Sturm’s sequence that the above expression is negative for $s$ in the given domain. Thus the proof is complete. $\square$
@@
Having characterised all pairs $(m,n)$ for which (1) holds, we now turn our attention to the non-solutions to this inequality. For convenience, we define the set
@@
For instance, we know that $\mathcal{I}^{2}_{n}=\emptyset$ for all $n$, and $(1/2,3/2,4/3)\in\mathcal{I}^{3}_{n}$ for $n\geqslant 5$. A trivial observation is if $\mathcal{I}_{n}^{m}\neq\emptyset$, then $\mathcal{I}_{n}^{m}$ is a nonempty open subset of $\mathcal{H}_{m}$. Consequently, $\mathcal{I}_{n}^{m}$ has positive (m1)(m-1)-dimensional measure with respect to the induced Lebesgue measure on this surface.
@@
$\lim_{x\to 0^{+}}S_{n}^{3}\left(x,\frac{1}{x},1\right)=\lim_{x\to+\infty}S_{n}^{3}\left(x,\frac{1}{x},1\right)=-1.$
@@
$x\in(0,\rho)\cup(\rho^{\prime},+\infty)\quad\Rightarrow\quad S_{n}^{3}\left(x,\frac{1}{x},1\right)<-\frac{7}{8}.$
@@
$S_{n}^{3}(x,y,z)<-\frac{3}{4},\quad\forall(x,y,z)\in B_{\infty}\left(\left(\rho,\frac{1}{\rho},1\right),\varepsilon_{f}\right).$ (6)
@@
$\Gamma\mathrel{\mathop{\ordinarycolon}}=\left\{\left(x,\frac{1}{x},1\right)\mathrel{\mathop{\ordinarycolon}}x\in[\rho,\rho^{\prime}]\right\}.$
@@
$\left\{B_{\infty}\left(\left(x,\frac{1}{x},1\right),\varepsilon_{x}\right)\mathrel{\mathop{\ordinarycolon}}\left(x,\frac{1}{x},1\right)\in\Gamma\setminus\{(1,1,1)\}\right\}\cup\{B_{\infty}((1,1,1),\varepsilon^{\prime})\}$
@@
$\left\{B_{\infty}\left(\left(x_{k},\frac{1}{x_{k}},1\right),\varepsilon_{k}\right)\right\}_{k=1}^{N}\subseteq\left\{B_{\infty}\left(\left(x,\frac{1}{x},1\right),\varepsilon_{x}\right)\right\}\cup\{B_{\infty}((1,1,1),\varepsilon^{\prime})\}$
@@
$|x-\tilde{\rho}|\leqslant\varepsilon_{f},\quad\left|y-\frac{1}{\tilde{\rho}}\right|\leqslant\varepsilon_{f},\quad|z-1|\leqslant\varepsilon_{f}.$
@@
$\left\|(x^{\prime},y^{\prime},z^{\prime})-\left(\rho,\frac{1}{\rho},1\right)\right\|_{\infty}=\left\|(x,y,z)-\left(\tilde{\rho},\frac{1}{\tilde{\rho}},1\right)\right\|_{\infty}\leqslant\varepsilon_{f}.$
@@
$S_{n}^{3}(x,y,z)=f(x)+f(y)+f(z)<f(x^{\prime})+f(y^{\prime})+f(z^{\prime})=S_{n}^{3}(x^{\prime},y^{\prime},z^{\prime})<0$
@@
for $x>0$. Assume to the contrary that $\mathcal{I}_{n}^{3}$ is unbounded. Then for every $c>0$, there exists $(x_{c},y_{c},z_{c})\in\mathcal{I}_{n}^{3}$ with $x_{c}<c$ where we assume without loss of generality that $x_{c}\leqslant y_{c}\leqslant z_{c}$. We record two elementary bounds.
@@
In addition to the above, it is natural to ask whether analogous structural properties hold for $\mathcal{I}_{n}^{m}$ when $m$ is arbitrary. Can one prove or disprove that $\mathcal{I}_{n}^{m}$ (i) stays uniformly away from $1$, (ii) is bounded, or (iii) forms an increasing family in $n$? Indeed, the answer to question (i) is negative according to Remark 4.2. We additionally note that the proof of Theorem 5.1 relies crucially on the fact that (1) holds for every $n\in\mathbf{N}$ when $m=2$, a feature that fails for other values of $m$, as shown in the previous sections. Because of this, we may likely think that the answer to question (ii) is negative as well. However, we do not know whether the converse “Theorem 5.2 $\Rightarrow$ Theorem 5.1” is true for any $m$. Moreover, we have not been able to construct an explicit counterexample yet. If it surprisingly turns out to be true, then it would require a new technique to prove. Thus this question, together with question (iii), remains open. Clarifying these properties of $\mathcal{I}_{n}^{m}$ for $m\geqslant 3$ would therefore be an interesting direction for further study, which we leave to the interested readers.